A 3-ft by 6-ft picnic table is placed on an outdoor patio measuring x ft by (x + 3) ft. Find a polynomial for the remaining open space on the patio (in square feet).

Answer by ankor@dixie-net.com(22740)

A 3-ft by 6-ft picnic table is placed on an outdoor patio measuring x ft by (x + 3) ft.
Find a polynomial for the remaining open space on the patio (in square feet).
:
3 * 6 = 18 sq/ft is the table
:
f(x) = x(x+3) - 18
f(x) = x^2 + 3x - 18